# Watch Your Language: Large Language Models and Content Moderation

### v1

@@ Methods
Improving rule-based prompting with behavior
@@
Parsing conversational context with LLMs
An ongoing challenge with all automated content moderation tools is in implementing conversational context. In our work, we found missing context to be a prevalent reason for false negatives and false positives in both tasks. Embedding context into toxicity detection has proved challenging in prior work (Pavlopoulos et al. 2020). However, most LLMs are already designed to be conversational in nature, and another potential area for future work might be identifying how best to leverage the conversational interface to simulate both conversation and moderation decisions.
@@
This work primarily focuses on text-based content moderation tasks, and specifically on comment moderation on discussion oriented platforms. As such, our results may not extend to other types of moderation, for example, those involving mixed modalities (i.e., images, videos, etc.) or those on social platforms with other affordances (i.e., real-time communication). Furthermore, our evaluation here is done on balanced datasets and thus our comment distributions do not reflect comment distributions found in real-world platforms. We also note that while our results are promising, the LLMs we tested are currently too expensive to immediately impact most content moderation contexts, and future research is needed to build models that can balance performance with cost.
@@ Conclusion
Overall, our results provide a tempered but optimistic view of introducing LLMs into content-moderation contexts. We hope that our work provides the groundwork for researchers and practitioners and charts a pathway forward for more work in studying how LLMs can shape and impact the future of content moderation.

### v2

@@ Discussion
Improving rule-based moderation with behavior and context
@@
Another promising direction to improve rule-based moderation is in incorporating conversational or world context into decision making. We observed how incorporating context can correct 35% of errors with little addition to the prompt, and future work can investigate incorporating other types of context (e.g., post details or images in a multimodal context) to aid in improving moderation outcomes.
@@
Future experiments in content moderation
Given that our LLMs performed better in a balanced context than in a real-world context, a promising area of future work would be to work collaboratively with moderators to see if LLMs might improve their existing moderation workflows. LLMs might help to steer moderators to potentially violative content, especially given our observation that  48% of false positives seem to violate community rules. Furthermore, moderators might have sufficient world-context to help steer LLMs through techniques like few-shot prompting (Brown et al. 2020) or through structure conversation (Argyle et al. 2023).
@@
This work primarily focuses on text-based content moderation tasks, and specifically on comment moderation on discussion oriented platforms. As such, our results may not extend to other types of moderation, for example, those involving mixed modalities (i.e., images, videos, etc.) or those on social platforms with other affordances (i.e., real-time communication). Furthermore, our evaluation here is primarily done on balanced datasets—our results with examining real-world distributions suggest that significantly more work needs to be done to evaluate such tools in practice. We also note that while are our results are promising, the LLMs we tested are too expensive to immediately impact most content moderation contexts, and future research is needed to build models that can balance performance with cost.
@@ Conclusion
Overall, our results provide a tempered but optimistic view of introducing LLMs into content-moderation contexts. We hope that our work provides the groundwork for researchers and practitioners and charts a pathway forward for more work in studying how LLMs can shape and impact the future of content moderation. We plan to release our code, prompts, and data at publication time to enable further research on automated content moderation with LLMs.
@@
Conversational Context Prompt
Consider the following comment tree:
<comment_1> : "<example_comment>"
<comment_2> : "<example_comment_2>"
Return a JSON object with five fields,
"would_moderate," that is
either "yes" or "no" depending
on if you would remove <comment_2>
from the subreddit, "rules" which are
the text of the rules being violated,
"rule_nums" which are a comma-separated
list of rules being violated, "rating"
which is a score from 1-5 on how
violative the comment is, and "explanation"
which provides a reason for your decision.
Cost and Performance of LLMs for Moderation
Subreddit
Metrics Toxic Task Rule Task
GPT-3
Time taken 3 days
Cost \$97
GPT-3.5
Time taken $<$1 day 5 days
Cost \$50 \$175
GPT-4
Time taken 2 days
Cost \$154
Gemini Pro
Time taken $<$1 day 5 days*
Cost N/A N/A
Table 10: Comparative cost and performance—We show each model’s approximate time taken and the cost of each experiment.
Table 10 shows a comparison between costs and overall “performance” (e.g., time taken to conduct the experiments) for each model we evaluate. We stress that such estimates, especially time taken, are challenging to reproduce as they are impacted by factors often out of our control, such as model load.